A credit manager rates each applicant for a car loan on a scale of 1-5 and then deterines the interest rate. Find the equation of a line in slope-intercept form that goes thru these points.
Credit Rating Interest Rate
1 24
2 20
3 16
4 12
5 8

Answer by cheesey(22)

first select two points to find the slope of the line. I chose (2, 20) (1, 24). Then the slope is (20-24)/(2-1) = -4
Using point-slope form: y - y* = m(x - x*), where y*, x* are one point given.
Thus y - 20 = -4(x - 2)
y - 20 = -4x + 8
y = -4x + 28 is one such equation of this line in y=mx + b form. Chose a different point and your "b" will change.